How much heat must be absorbed by water of mass ? = 300 ? at ? = 20°? to take it to the gas state at 100°?? (The heat of vaporization is 540 ??? /? and the specific heat of water is 1.00 ??? /?? )
Answer #1

Q,+ Q2 2 PE Q = m.c.ST Q = 300 X / X (100-20) Q = 300×80 cal = 24000 cal Q2 = m.. Q2 =300x540 Q2 cal 162ooo cal 2 Q = Q 2 yooPlease UpVote and let
me know if you face any problem, confusion or see any error.
I will get back to you asap.